NH3 is a weak base (Kb = 1.8 × 10–5) and so the salt NH4Cl acts as a weak acid. What is the pH of a solution that is 0.044 M in NH4Cl at 25 °C?
Answer #1

This will happen:

NH$Cl = NH4+ + Cl-.

NH4+ + H2O <-> NH3 + H3O+

Ka = [NH3 ][H3O+]/[NH4+]

Ka = Kw/Kb = (10^-14)/(1.8*10^-5) = 5.55*10^-10

[NH3 ]= [H3O+]= x

[NH4+] = M-x = 0.044-x

Ka = [NH3 ][H3O+]/[NH4+]

5.55*10^-10 = x*x/(0.044-x)

x = 4.94*10^-6

H+ =  4.94*10^-6

pH = -log( 4.94*10^-6) =5.30627